# LLMs for Literature Review: Are we there yet?

### v1

@@ 7 Conclusions & Answering: Are We There Yet?
This work discusses, establishes and evaluates a pipeline to help people write literature reviews. To explore this problem setting we have proposed and implemented a rolling evaluation procedure and collected several recent datasets which we will make available to the community. We show that LLMs show significant potential for writing literature reviews when the task is decomposed into these smaller and simpler sub-tasks that are within the reach of LLMs, namely through the use of LLM generated keword search and embedding based search for relevant prior work. Notably, our experiments indicate that attribution based on citing extracted content from source material improves LLM re-ranking results. The most powerful LLMs evaluated in our studies exhibit both extremely promising paper re-ranking abilities as well as promising literature review generation results. Importantly, LLM hallucinations can be substantially reduced using our proposed plan based prompting and retrieval augmented generation techniques. Our evaluation also reveals clear challenges: 1) retrieving all relevant papers consistent with a given human generated literature review will requires new querying strategies; 2) hallucinations can be significantly reduced using plan-based prompting, but our approach does not completely eliminate hallucinations. Moreover, based on the proposed retrieval and generation pipeline, we build a full working demo (see Figure 15 of our Appendix for a screenshot), which we will release to the community. We hope that authors can use this demo to better understand how these techniques and future alternatives can be most helpful for literature review generation assistance.
@@
Limitations and Future Work Because of the low coverage for retrieval, we evaluate different components independently. During generation, this strategy assumes that we already have filtered relevant papers corresponding to the main paper. In the future, we would like to improve the search for relevant work using embedding-based models to get better coverage and, thus, the ability to evaluate the system end-to-end. Our retrieval component currently suffers from surface-level information about the papers, while in practice, authors frame search keywords based on information (such as underlying datasets) which might not be present in the abstract. Another issue stems from the retrieval evaluation setup based on coverage related to the ground truth papers, where the authors might have different biases. We also acknowledge that including more of the paper, like introduction and methodology, might help improve the set of initial candidate papers; however, using additional sections could inadvertently allow the model to detect explicit citations or references, essentially “cheating” by hinting at specific papers to retrieve. By focusing on the abstract alone, we maintain a more controlled setup that reflects a realistic, early-stage research scenario. It is important to highlight that while we operate in this abstract-only setup, our pipeline is designed to be flexible and interactive. As the paper matures and more content becomes available, researchers can provide additional context to the pipeline to improve retrieval accuracy. This adaptability ensures that our approach remains relevant and effective throughout different stages of the research process, allowing for incremental refinement of related work as drafts evolve over time.
@@ Appendix A Ethics Statement
The potential for LLM and other NLP technology to help in scientific writing has led to the emergence of systems such as Explainpaper which helps researchers understand the contents of the paper and Writefull [^12] https://www.explainpaper.com/, https://x.writefull.com/ for title and abstract generation. Scite [^13] https://scite.ai/ helps find appropriate resources and references while writing research papers. The usage of LLMs by researchers is so evident that some conferences (like ICLR) collect statistics from authors about their usage of these tools for retrieval (or discovery) or paraphrasing the related work section and provide LLM-specific guidelines to authors. [^14] ICLR’24 Large Language Models guidelines https://iclr.cc/Conferences/2024/CallForPapers While writing assistant technology could have great promise as an aide to scientists, we think their use should be disclosed to the reader. As such assistants become more powerful they might be abused in certain contexts, for example where students are supposed to create a literature review as a part of their learning process. The use of such tools might also be problematic as authors of scientific work should read the articles that they cite and heavy reliance on such tools could lead to short term gains at the cost of a deeper understanding of a subject over the longer term. Any commercially deployed or systems actually used by authors should also contain appropriate mechanisms to detect if words have been copied exactly from the source material, and provide that content in a quoted style. Additionally, the rolling evaluations which we present here does not involved training LLMs on arXiv papers. This mitigates concerns regarding the copyright status of arXiv papers and their use for LLM training.
@@ Appendix B New Datasets
While there are datasets available for different tasks in academic literature (see Table 7), we use the Multi-XScience dataset  (Lu et al., 2020) for our experiments. Recent work  (Chen et al., 2021b; Funkquist et al., 2022) also focuses on related work generation and provides a similar dataset. As part of this work, we release two corpora: 1. We extend the Multi-XScience corpus to include full-text of research papers and 2. We create a new test corpus RollingEval-Aug consisting of recent (August 2023) arXiv papers (with full content).
@@
BigSurvey-MDS (Liu et al., 2023) Survey Introduction
@@
SciXGen (Chen et al., 2021a) Context aware text generation
@@
Multi-XScience full text We create these datasets based on the latest release (2023-09-12) of the S2ORC corpus [^15] Dataset available at http://api.semanticscholar.org/datasets/v1/  (Lo et al., 2020) available at the Semantic Scholar Open Data Platform  (Kinney et al., 2023). The S2 Platform provides access to multiple datasets including papers metadata, authors, S2AG (Semantic Scholar Academic Graph), paper embeddings, etc. While the ‘Papers’ dataset consists of 200M+ metadata records, S2ORC consists of 11+M full-text publicly available records with annotations chunked into 30 files ($\sim$215G compressed json) where research documents are linked with arXiv and Microsoft Academic Graph (MAG)  (Sinha et al., 2015) IDs, when available. This corpus provides full text of the research papers (parsed using a complex pipeline consisting of multiple LaTeX and PDF parsers such as GROBID  (Lopez, 2023) and in-house parsers. [^16] https://github.com/allenai/papermage). The full text is also aligned with annotation spans (character level on the full text), which identify sections, paragraphs, and other useful information. It also includes spans for citation mentions, and the matching semantic corpus-based ID for bibliographical entries making it easier to align with references compared to other academic datasets such as LoRaLay  (Nguyen et al., 2023), UnarXive  (Saier & Färber, 2020; Saier et al., 2023), etc. or relying on citation graphs like OpenAlex  (Priem et al., 2022), next-generation PDF parsers  (Blecher et al., 2023) or other HTML webpages. [^17] https://ar5iv.labs.arxiv.org/ and https://www.arxiv-vanity.com/ For the Multi-XScience, we obtain the full text of papers for 85% of records from the S2ORC data using the span annotations from the corpus aligned with citation information.
@@
RollingEval datasets Llama 2 was publicly released on 18th July 2023 and GPT-4 on 14 March 2023. Both provide limited information about their training corpus where academic texts in the Multi-XScience may or may not have been part of their training data. To avoid overlap with the training data of these LLMs, we process a new dataset using papers posted after their release date. To do so, we first filter the papers published in August 2023 from S2ORC that contain an arXiv ID resulting in $\sim$15k papers. S2ORC does not provide the publication date of the papers directly, so we use regex ‘2308’ on the arXiv ID to extract papers posted in 08’23. We then use section annotations to get the section names and match using synonyms (‘Related Work, Literature Review, Background’) to extract section spans. We take the rest of the text as conditioning context except the related work section which results in $\sim$4.7k documents. Using the citation annotations, we extract the full text of cited papers from the S2ORC corpus again using corpus ID. Similar to Multi-XScience, we use paragraph annotations to create a dataset for the latest papers ($\sim$6.2k rows). We create a subset of 1,000 examples (RollingEval-Aug) where we have the content of all the cited papers. Average length of related work summary is 95 words while the average length of abstracts is 195. On an average, we have 2 citations per example which makes the dataset comparable to the original Multi-XScience dataset.
@@ Appendix C Other Generation Experiments
In parallel, we also fine-tune Llama 2 models on the train set with the original shorter context, but they are very sensitive to hyperparameter configuration. When we instruct-finetune Llama 2 7B, it initially produces code. We find a slight improvement when fine-tuning the Llama 2 7B model for 30k steps with an LR of 5e-6 over 0-shot model (see Table 8), but it quickly overfits as we increase the LR or the number of steps. We leave hyperparameter optimization, the fine-tuning of larger models with RoPE scaling and plan-based generation for future work.
@@
Code LLMs We evaluate the performance of code-generating LLMs to write related-work section which requires more formal and structured language. Since Code LLMs are pre-trained on text they might offer the best of both worlds. However, we observe that for our task, the models produce bibtex and Python code with relevant comments as part of the generated outputs. As shown in Table 10, CodeLlama (34B Instruct) is good at following instructions and at generating natural language (ROUGE2 of 5.8 and 5.02 on Multi-XScience and RollingEval-Aug dataset). With a plan, CodeLlama even surpasses vanilla 0-shot Llama 2 70B (Table 4).
@@
Table 10: 0-shot results using code-based models on Multi-XScience dataset. CodeLlama performs reasonably well in generating natural language compared to the other code based counterparts.
@@ Appendix D More implementation details
D.1 Generation Implementation
@@
D.2 Demo implementation
We build our system using Gradio (Abid et al., 2019), which provides a nice interface to quickly and efficiently build system demos. Our user interface is also available at HuggingFace Space [^24] https://huggingface.co/spaces/.
We query the Semantic Scholar API available to search for the relevant papers. Specifically, we use the Academic Graph [^25] https://api.semanticscholar.org/api-docs/graph and Recommendations [^26] https://api.semanticscholar.org/api-docs/recommendations API endpoint. We use OpenAI API to generate results for LLM using GPT-3.5-turbo and GPT-4 model. At the same time, our modular pipeline allows using any LLM (proprietary or open-sourced) for different components. We also allow the end-user to sort the retrieved papers by relevance (default S2 results), citation count, or year. More details about the demo system can be found in our system paper.
D.3 SPECTER Implementation
We build an index of 150M SPECTER2 embeddings that we can use as an alternative to both a search engine and a prompting-based ranking module. Figure 6 shows our pipeline for creating the index. Specifically, the SPECTER2 database comes with 908 json.gz files containing compressed embeddings. For each json.gz file, we construct a FAISS index that we can query for the nearest neighbors of a given query embedding. We perform index construction in a multi-threaded manner to speed up the process. Upon constructing a FAISS index for all the json.gz files, we iterate over each query paper, search for top 100 relevant papers using the SPECTER embeddings in each FAISS index, and then finally merge the results to get the top 1000 papers for each query paper.
@@ D.3 SPECTER Implementation
Table 11: Abstracts of papers which are reference citations in Tables 12 and 13.
@@
Table 12: We show outputs of different models corresponding to the same plan and prompt as paragraph of our related work (by GPT-4 Plan). Llama 2 70B 0-shot has the tendency to hallucinate citation (in red) which is remedied in plan based generation. Abstract of reference paper shown in Table 11 while the main abstract is a slight modified version of our current abstract.
@@
In this section, we discuss the related work in the field of scientific literature review generation using large language models (LLMs). Recent works have explored the use of LLMs for multi-document summarization tasks, such as generating the related work section of a paper based on its abstract and the articles it references (@cite_1). These works have shown that LLMs can produce high-quality summaries, but may suffer from the lack of relevant information and context (@cite_2). To address this issue, we propose a novel strategy for literature review generation that involves generating a plan for the review before generating the actual text. This approach has been shown to improve the quality of generated literature reviews over vanilla zero-shot generation (@cite_3). Additionally, we create a new test corpus consisting of recent arXiv papers to ensure that our zero-shot experiments do not suffer from test set contamination. Our work builds upon the recent advances in LLMs and their ability to store, combine, and reason about scientific knowledge (@cite_3). We believe that our approach has the potential to revolutionize the way scientific literature is reviewed and organized, and we open source our model for the benefit of the scientific community. In summary, our related work demonstrates the potential of LLMs for scientific literature review generation, and highlights the importance of context and information in producing high-quality summaries. Our proposed approach aims to address these limitations and provide a more effective way of generating literature reviews using LLMs.
Table 13: We show outputs of different models corresponding to the reference cited abstracts and plan mentioned in Table 11 while the main abstract is a slight modified version of our current abstract. In this example, though, we have all the citations covered by all the models, we can see GPT-4 (Plan) output to be concise and closely following the plan.
@@
Figure 10: Prompt used for Plan based generation.
Figure 11: Prompt used when plan is learned during generation.
Figure 12: Prompt used for Sentence by sentence generation.
@@
Figure 15: Anonymized system interface on Huggingface Space. Our system works on the Retrieval Augmented Generation (RAG) principle to generate the literature review grounded in retrieved relevant papers. User needs to provide the abstract in the textbox (in purple) and press send to get the generated related work (in red). First, the abstract is summarized into keywords, which are used to query a search engine. Retrieved results are re-ranked (in blue) using an LLM, which is then used as context to generate the related work. Users could also provide a sentence plan (in green) according to their preference to generate a concise, readily usable literature review.

### v2

@@ 7 Conclusions & Answering: Are We There Yet?
This work discusses, establishes and evaluates a pipeline to help people write literature reviews. We first identified some challenges of evaluating such systems when LLMs are constantly updated based on training on new data, which may contain recent papers found online. To address these issues, we propose and implement a rolling evaluation procedure that focuses on recent arXiv papers, and we collect several evaluation datasets in this manner.
@@
Our experiments show that LLMs have significant potential for writing literature reviews, especially when the task is decomposed into these smaller and simpler sub-tasks that are within reach of LLMs, namely through the use of LLM-generated keyword search and embedding-based search for relevant prior work. Notably, our experiments indicate that both debate prompting and debate arguments that use attribution based on citing extracted content from source material improve LLM re-ranking results. The most powerful LLMs evaluated in our studies exhibit extremely promising paper re-ranking abilities as well as promising literature review generation results. Importantly, LLM hallucinations can be substantially reduced using our proposed plan-based prompting and retrieval augmented generation techniques.
Our evaluation also reveals clear challenges: 1) retrieving all relevant papers consistent with a given human-generated literature review will require new querying strategies; 2) hallucinations can be significantly reduced using plan-based prompting, but our approach does not completely eliminate hallucinations.
So, are we there yet? Not quite—but we are getting closer. The landscape of AI-assisted research exploded in 2025, with tools like DeepResearch (OpenAI, 2025), AI Co-Scientist (Gottweis et al., 2025), and ScholarQA (AllenAI, 2025) demonstrating remarkable improvements in literature review generation, citation accuracy, and retrieval strategies. Moreover, to accompany our scientific work here, and our older work on this theme (Agarwal et al., 2024), we have build a full working demo based on our proposed retrieval and generation pipeline (see Figure 16 in the Appendix for a screenshot), which we will release to the community. We hope that authors can use this demonstration system to better understand how these techniques—and future alternatives—can be most helpful for assistance in literature review generation.
@@
Limitations and Future Work. Because of the low coverage for retrieval, we evaluate different components independently. During generation, this strategy assumes that we already have filtered relevant papers corresponding to the main paper. In the future, we would like to improve the search for relevant work using embedding-based models to get better coverage and, thus, the ability to evaluate the system end-to-end. Our retrieval component currently suffers from surface-level information about the papers, while in practice, authors frame search keywords based on information (such as underlying datasets) that might not be present in the abstract. Another issue stems from the retrieval evaluation setup based on coverage related to the ground truth papers, where the authors might have different biases. We also acknowledge that including more of the paper, like introduction and methodology, might help improve the set of initial candidate papers; however, using additional sections could inadvertently allow the model to detect explicit citations or references, essentially “cheating” by using these as hints to retrieve specific papers. By focusing on the abstract alone, we maintain a more controlled setup that reflects a realistic, early-stage research scenario. It is important to highlight that while we operate in this abstract-only setup, our pipeline is designed to be flexible and interactive. As the paper matures and more content becomes available, researchers can provide additional context to the pipeline to improve retrieval accuracy. This adaptability ensures that our approach remains relevant and effective throughout different stages of the research process, allowing for incremental refinement of related work as drafts evolve.
@@ Appendix A Ethics Statement
The rapid advancements in LLMs and NLP technologies for scientific writing have led to the emergence of increasingly powerful systems such as DeepResearch, AI Co-Scientist, and ScholarQA. These tools extend beyond earlier systems like Explainpaper and Writefull [^12] https://www.explainpaper.com/, https://x.writefull.com/, which assist in paper comprehension and abstract generation, and Scite [^13] https://scite.ai/, which helps with citation discovery. As AI-powered tools become more deeply integrated into the scientific workflow, ethical considerations around their use continue to evolve. Many conferences, such as ICLR, have begun collecting statistics on authors’ usage of LLMs for literature review generation and paraphrasing, and have issued guidelines on responsible usage. [^14] ICLR’24 Large Language Models guidelines https://iclr.cc/Conferences/2024/CallForPapers While writing assistant technology could have great promise as an aide to scientists, we think their use should be disclosed to the reader. As such assistants become more powerful, they might be abused in certain contexts, for example, where students are supposed to create a literature review as a part of their learning process. The use of such tools might also be problematic as authors of scientific work should read the articles that they cite, and heavy reliance on such tools could lead to short-term gains at the cost of a deeper understanding of a subject over the longer term. Any commercially deployed systems authors use should also contain appropriate mechanisms to detect if words have been copied exactly from the source material and provide that content in a quoted style. Additionally, as newer tools like DeepResearch, AI Co-Scientist, and ScholarQA continue to improve, it is crucial to assess their long-term impact on scientific research. The use of these tools should complement, rather than replace, human expertise in literature analysis. Finally, the rolling evaluations we present here do not involve training LLMs on arXiv papers. This mitigates concerns regarding the copyright status of arXiv papers and their use for LLM training.
@@ Appendix B New Datasets
While there are datasets available for different tasks in academic literature (see Table 7), we use the Multi-XScience dataset  (Lu et al., 2020) for our experiments. Recent work  (Chen et al., 2021b; Funkquist et al., 2022) also focuses on related work generation and provides a similar dataset. As part of this work, we release two corpora: 1. We extend the Multi-XScience corpus to include the full text of research papers, and 2. We create a new test corpus, RollingEval-Aug, consisting of recent (August 2023) arXiv papers (with full content).
@@
BigSurvey-MDS (Liu et al., 2023a) Survey Introduction
@@
SciXGen (Chen et al., 2021a) Context-aware text generation
@@
Multi-XScience full text We create these datasets based on the latest release (2023-09-12) of the S2ORC corpus [^15] Dataset available at http://api.semanticscholar.org/datasets/v1/  (Lo et al., 2020) available at the Semantic Scholar Open Data Platform  (Kinney et al., 2023). The S2 Platform provides access to multiple datasets, including paper metadata, authors, S2AG (Semantic Scholar Academic Graph), paper embeddings, etc. While the ‘Papers’ dataset consists of 200M+ metadata records, S2ORC consists of 11+M full-text publicly available records with annotations chunked into 30 files ($\sim$215G compressed json) where research documents are linked with arXiv and Microsoft Academic Graph (MAG)  (Sinha et al., 2015) IDs, when available. This corpus provides full text of the research papers (parsed using a complex pipeline consisting of multiple LaTeX and PDF parsers such as GROBID  (Lopez, 2023) and in-house parsers. [^16] https://github.com/allenai/papermage). The full text is also aligned with annotation spans (character level on the full text), which identify sections, paragraphs, and other useful information. It also includes spans for citation mentions and the matching semantic corpus-based ID for bibliographical entries, making it easier to align with references compared to other academic datasets such as LoRaLay  (Nguyen et al., 2023), UnarXive  (Saier & Färber, 2020; Saier et al., 2023), etc. or relying on citation graphs like OpenAlex  (Priem et al., 2022), next-generation PDF parsers  (Blecher et al., 2023) or other HTML webpages. [^17] https://ar5iv.labs.arxiv.org/ and https://www.arxiv-vanity.com/ For the Multi-XScience, we obtain the full text of papers for 85% of records from the S2ORC data using the span annotations from the corpus aligned with citation information.
@@
RollingEval datasets Llama 2 was publicly released on 18th July 2023 and GPT-4 on 14 March 2023. Both provide limited information about their training corpus, and academic texts in the Multi-XScience may or may not have been part of their training data. To avoid overlap with the training data of these LLMs, we process a new dataset using papers posted after their release date. To do so, we first filter the papers published in August 2023 from S2ORC that contain an arXiv ID, resulting in $\sim$15k papers. S2ORC does not provide the publication date of the papers directly, so we use regex ‘2308’ on the arXiv ID to extract papers posted in 08’23. We then use section annotations to get the section names and match using synonyms (‘Related Work, Literature Review, Background’) to extract section spans. We take the rest of the text as conditioning context except the related work section which results in $\sim$4.7k documents. Using the citation annotations, we extract the full text of cited papers from the S2ORC corpus again using corpus ID. Similar to Multi-XScience, we use paragraph annotations to create a dataset for the latest papers ($\sim$6.2k rows). We create a subset of 1,000 examples (RollingEval-Aug) where we have the content of all the cited papers. The average length of a related work summary is 95 words, while the average length of abstracts is 195. On average, we have 2 citations per example, which makes the dataset comparable to the original Multi-XScience dataset.
@@ Appendix C Other Generation Experiments
In parallel, we also fine-tune Llama 2 models on the train set with the original shorter context, but they are very sensitive to hyperparameter configuration. When we instruct-finetune Llama 2 7B, it initially produces code. We find a slight improvement when fine-tuning the Llama 2 7B model for 30k steps with an LR of 5e-6 over 0-shot model (see Table 8), but it quickly overfits as we increase the LR or the number of steps. We leave hyperparameter optimization, fine-tuning larger models with RoPE scaling and plan-based generation for future work.
@@
Code LLMs We evaluate the performance of code-generating LLMs to write related-work sections requiring more formal and structured language. Since Code LLMs are pre-trained on text they might offer the best of both worlds. However, we observe that for our task, the models produce bibtex and Python code with relevant comments as part of the generated outputs. As shown in Table 10, CodeLlama (34B Instruct) is good at following instructions and at generating natural language (ROUGE2 of 5.8 and 5.02 on Multi-XScience and RollingEval-Aug dataset). With a plan, CodeLlama even surpasses vanilla 0-shot Llama 2 70B (Table 4).
@@
Table 10: 0-shot results using code-based models on Multi-XScience dataset. CodeLlama performs reasonably well in generating natural language compared to the other code-based counterparts.
@@ Appendix D More implementation details
D.1 Normalized Recall v/s Standard Recall: A Worked-out Example
Consider a query paper with the following statistics:
|Ground $|\text{Ground Truth}|=n_{\text{gt}}=84$
$|\text{Retrieved}|=100$
Relevant Retrieved papers=|RetrievedGround $\text{Relevant Retrieved papers}=|\text{Retrieved}\cap\text{Ground Truth}|=c=10$
[…]
Using these values, we compute the metrics at $k=40$:
[…]
This example illustrates how Normalized Recall@k differs from standard recall. Instead of being limited by the total number of ground truth citations, it evaluates how well the method ranks the retrievable relevant papers. In this case, despite a low precision, the normalized recall is relatively high, indicating that the method effectively ranks the relevant papers it does retrieve.
D.2 Generation Implementation
@@
D.3 Demo implementation
We build our system using the ReactJS framework, which provides a nice interface to build system demos quickly and efficiently. More details about the demo implementation can be found in our system paper Agarwal et al. (2024) and the project page.
We query the Semantic Scholar API available to search for the relevant papers. Specifically, we use the Academic Graph [^24] https://api.semanticscholar.org/api-docs/graph and Recommendations [^25] https://api.semanticscholar.org/api-docs/recommendations API endpoint. We use OpenAI API to generate results for LLM using GPT-3.5-turbo and GPT-4 models. At the same time, our modular pipeline allows using any LLM (proprietary or open-sourced) for different components. We also allow the end-user to sort the retrieved papers by relevance (default S2 results), citation count, or year. More details about the demo system can be found in our system paper.
D.4 SPECTER Implementation
We build an index of 150M SPECTER2 embeddings that we can use as an alternative to both a search engine and a prompting-based ranking module. Figure 6 shows our pipeline for creating the index. Specifically, the SPECTER2 database comes with 908 json.gz files containing compressed embeddings. For each json.gz file, we construct a FAISS index that we can query for the nearest neighbors of a given query embedding. We perform index construction in a multi-threaded manner to speed up the process. Upon constructing a FAISS index for all the json.gz files, we iterate over each query paper, search for the top 100 relevant papers using the SPECTER embeddings in each FAISS index, and then finally merge the results to get the top 1000 papers for each query paper.
D.5 Comparative analysis of the computational costs
We compare the costs of different LLMs for both stages in Table 11.
Ranking: We explore two types of LLM-based reranking mechanisms: permutation and debate ranking. For $n$ query papers (=500 for our RollingEval datasets) and top-$k$ candidates retrieved from S2 per query paper ($k$=100 in our experiments), permutation ranking would require $n$ API calls, whereas debate ranking would require $n*k$ API calls. Debate ranking needs more API calls as it involves one additional API call per candidate paper to generate the citation probability score and reasoning. Therefore, there are $k$ additional API calls per query paper compared to permutation ranking, where we prompt the LLM to directly rank relevance for all the candidate papers. We refer the reader to Figure 15for The exact prompt used for debate ranking.
Generation: There was only one request per query abstract in the RollingEval dataset, so 500 requests in total for each experiment (as $n=500$ in RollingEval). The table below summarizes the API analysis for the two stages of the pipeline for the RollingEval experiments.
Experiment Method Requests Tokens Cost
Ranking GPT-4 Permutation Reranking 500 $\sim$20M input + $\sim$0.25M output tokens \$50
Llama-3.1 Debate Ranking (w/o attribution) 500 x 100 $\sim$33M input + $\sim$0.25M output tokens \$0
Llama-3.1 Debate Ranking (w/ attribution) 500 x 100 $\sim$33M input + $\sim$15M output tokens \$0
Generation Llama 2 70B (using Anyscale Endpoint) 500 $\sim$0.75M input + $\sim$0.15M output tokens \$3.84
GPT-3.5-turbo 500 $\sim$0.75M input + $\sim$0.15M output tokens \$4.2
GPT-4 500 $\sim$0.75M input + $\sim$0.15M output tokens \$22
GPT-4 (Plan) 500 $\sim$0.75M input + $\sim$0.15M output tokens \$25
Table 11: Computational costs for different experiments on the RollingEval dataset. Costs for generation experiments on the Multi-XScience are approximately 10 times that of the RollingEval dataset.
@@ D.5 Comparative analysis of the computational costs
Table 12: Abstracts of papers which are reference citations in Tables 13 and 14.
@@
Table 13: We show outputs of different models corresponding to the same plan and prompt as paragraph of our related work (by GPT-4 Plan). Llama 2 70B 0-shot has the tendency to hallucinate citation (in red) which is remedied in plan based generation. Abstract of reference paper shown in Table 12 while the main abstract is a slight modified version of our current abstract. We also find that the 0-shot results are more verbose, copying the exact text verbatim, while the generated literature review using plans are more representative of the way humans write literature review.
@@
In this section, we discuss the related work in the field of scientific literature review generation using large language models (LLMs). Recent works have explored the use of LLMs for multi-document summarization tasks, such as generating the related work section of a paper based on its abstract and the articles it references (@cite_1). These works have shown that LLMs can produce high-quality summaries, but may suffer from the lack of relevant information and context (@cite_2). To address this issue, we propose a novel strategy for literature review generation that involves generating a plan for the review before generating the actual text. This approach has been shown to improve the quality of generated literature reviews over vanilla zero-shot generation (@cite_3). Additionally, we create a new test corpus consisting of recent arXiv papers to ensure that our zero-shot experiments do not suffer from test set contamination. Our work builds upon the recent advances in LLMs and their ability to store, combine, and reason about scientific knowledge (@cite_3). We believe that our approach has the potential to revolutionize the way scientific literature is reviewed and organized, and we open-source our model for the benefit of the scientific community. In summary, our related work demonstrates the potential of LLMs for scientific literature review generation and highlights the importance of context and information in producing high-quality summaries. Our proposed approach aims to address these limitations and provide a more effective way of generating literature reviews using LLMs.
Table 14: We show outputs of different models corresponding to the reference cited abstracts and plan mentioned in Table 12 while the main abstract is a slightly modified version of our current abstract. In this example, though, we have all the citations covered by all the models, we can see GPT-4 (Plan) output to be concise and closely following the plan. Moreover, GPT-3.5 and Llama 0-shot outputs are excessively elaborate, making them unsuitable for inclusion in the literature review. Our findings indicate that while LLMs can help in certain aspects, the final output still heavily relies on inputs from a human researcher.
@@
Figure 10: Prompt used for plan-based generation. Underlined text shows the variation compared to the vanilla 0-shot prompting, where the user provides a structure of the expected paragraph.
Figure 11: Prompt used when the plan is learned during generation. The model first generates a plan of sentences and citations which it would then condition upon to generate the final related work text, which can be considered as an extension of CoT style thinking step by step.
Figure 12: Prompt used for sentence-by-sentence generation. In this scenario, we prompt the model to generate one sentence for each citation individually.
@@
You are a helpful research assistant who is helping with literature review of a research idea. Your task is to rank some papers based on their relevance to the query abstract.
## Instruction:
Given the query abstract:
<query_abstract>{query_abstract}</query_abstract>
Given the candidate reference paper abstract:
<candidate_paper_abstracts>{reference_papers}</candidate_paper_abstracts>
* Given the abstract of the candidate reference papers, provide me with a number between 0 and 100 (upto two decimal places) that is proportional to the probability of a paper with the given query abstract including the candidate reference paper in its literature review.
* In addition to the probability, give me arguments for and against including this paper in the literature review.
* You must enclose your arguments for including the paper within <arguments_for> and </arguments_for> tags.
* You must enclose your arguments for including the paper within <arguments_against> and </arguments_against> tags.
* Extract relevant sentences from the candidate paper abstract to support your arguments.
* Put the extracted sentences in quotes.
* You can use the information in other candidate papers when generating the arguments for a candidate paper.
* You must enclose your score within <probability> and </probability> tags.
* Generate the arguments first then the probability score.
* Generate arguments and probabitlity for each paper separately.
* Do not generate anything else apart from the probability and the arguments.
* Follow this process even if a candidate paper happens to be identical or near-perfect match to the query abstract.
### Response Format for each paper:
<arguments_for>
[Paper ID]: [Reason for including the paper]
Extracted Sentences: "Sentence 1", "Sentence 2",
</arguments_for>
<arguments_against>
[Paper ID]: [Reason for not including the paper]
Extracted Sentences: "Sentence 1", "Sentence 2",
</arguments_against>
<probability>
[Paper ID]: [Final Probability Score Based on the Arguments]
</probability>
### Your Response:
Figure 15: Prompt used for Debate Ranking.
Figure 16: LitLLM interface Agarwal et al. (2024). Our system works on the Retrieval Augmented Generation (RAG) principle to generate the literature review grounded in retrieved relevant papers. The user needs to provide the abstract in the textbox (in purple) and press send to get the generated related work (in red). First, the abstract is summarized into keywords, which are used to query a search engine. Retrieved results are re-ranked (in blue) using an LLM, which is then used as context to generate the related work. Users could also provide a sentence plan (in green) according to their preference to generate a concise, readily usable literature review.